1.

2.

3.

If there is a negative correlation between X and Y then the regression equation, γ= bX + a will have Ob>o a>0 a<0

Which of the following sets of sample data would produce the largest value for an independent-measures t statistic. Assume that n = 10 for all samples. Note: You should not need to do any serious calculations to answer this question. First sample: M 30 and SS 10, Second Sample: M 35 and SS 10 ● First sample : M = 30 and SS = 50, Second Sample: M = 35 and SS = 50 O First sample: M 30 and SS 10, Second Sample: M 50 and SS 10 O First sample: M 30 and SS 50, Second Sample: M50 and SS 50
A college professor reports that students who finish exams early tend to get better grades than students who hold on to exams until the last possible moment. The correlation between exam score and amount of time spent on the exam is an example of a positive correlation a negative correlation. a correlation near zero a correlation near one.
If there is a negative correlation between X and Y then the regression equation, γ= bX + a will have Ob>o a>0 a
Answer #1

1)

Given

Test statistic t = (M1-M2)/sqrt(s^2(1/n1+1/n2))

Options

a) t = (30-35) / sqrt (10^2 (1/10 + 1/10) ) = -1.118

b) t = -0.01

c) t = -1

d) t = -0.8944

Option b would produce largest t-statistic.

2)

Negative correlation.

3)

b < 0